[figure: 9.7 cm]

EFFECT OF STRUCTURAL DISORDERING ON MAGNETIC PROPERTIES OF STOICHIOMETRIC Ni_{2}MnGa ALLOY FILMS

Y.V. Kudryavtsev^{a}, J. Dubowik^{b},Y.P. Lee^{c} aInstitute of Metal Physics, 36 Vernadsky Ave., 03142 Kiev, Ukraine bInstitute of Molecular Physics, Polish Academy of Siences, M. Smoluchowskiego 17, 61-179, Poznań, Poland

cDepartment of Physics, Hanyang University, 17 Haendang-Dong, Sungdong-Ku, Seoul 133-791 Korea

Ni_{2}MnGa Heusler alloy is currently extiting considerable interest due to giant ferromagnetic shape memory effect. This property makes thin films of Ni_{2}MnGa promising candidates for microelectromechanical systems. We report our results concerning structure, magnetic and transport properties of nearly stoichiometric Ni_{2}MnGa films prepared by flash-evaporation. Two distinct substrate temperatures (150 and 720 K) were chosen to fabricate films with different long-range order. We show for the first time that ordered films deposited at 720 K exhibit the same magnetic and transport properties as those of crystalline bulk Ni_{2}MnGa shape memory alloys with the martensitic transformation at 200 K, Curie temperature of 380 K and saturation magnetization at room temperature of about 300 G. Disordered films deposited at 150 K are shown to behave as the highly resistive Pauli paramagnets. Annealing of the disordered Ni_{2}MnGa films restores the ordered structure with almost full recovery of the magnetic and transport properties of the ordered Ni_{2}MnGa films.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : Yu. V. Kudryavtsev

Address for correspondence : Institute of Metal Physics, 36 Vernadsky Ave., 03142 Kiev, Ukraine

Email address : kudr@imp.kiev.ua